OpenStudy (anonymous):

solve for x 2 = x + e^x

OpenStudy (blockcolder):

The only thing I'm certain about here is that the solution lies between 0 and 1.

OpenStudy (anonymous):

No dp stipulated, is there an exact solution?

jhonyy9 (jhonyy9):

what do you think from this - make logarithm naturale from both sides ,like lne^2 =lne^x +lne^x - so if this is right ?

OpenStudy (anonymous):

I don't follow that....

jhonyy9 (jhonyy9):

so than e^2=e^2x than 2=2x x=2/2 x=1

OpenStudy (blockcolder):

But 1+e^1 is not 2.

OpenStudy (precal):